OpenStudy (anonymous):

how do you solve 2lnx+3ln2=5 ?

myininaya (myininaya):

2lnx=5-3ln2 lnx=1/2*(5-3ln2) x=e^(.5*(5-3ln2))

OpenStudy (anonymous):

ln[x²·2³]= 5 --> 8x²= e^5 --> x²= 1/8· e^5 and x>0 --> x= √2/4 ·e^(5/2)~ 4.3071....

OpenStudy (anonymous):

hello mininaya!

myininaya (myininaya):

hey